# The Winner of the NFL Draft is Not Necessarily Cursed

### v1

@@ 1.2 Traditional draft value curves
The fitted trade market and Jimmy Johnson curves are quite similar. This is not a coincidence: even today, some teams still trade using the Jimmy Johnson chart. The trade market curve is much steeper than expected performance and surplus value curves. In other words, teams aren’t valuing draft picks by their expected value. The implication is that general managers are trading suboptimally: they don’t trade down often enough, and they often overpay when they trade up. According to surplus value, even teams at the very top of the draft should trade down, particularly towards the middle of the first round. This result is counterintuitive because very top picks are so highly regarded by general managers and football fans.
1.3 Research hypothesis
Massey and Thaler, (2013) pioneered the traditional analytical approach to calculating NFL draft curves, which are expected value curves. They had two main findings. They discovered the loser’s curse: according to expected surplus value, top picks aren’t as valuable as later picks. Also, the way general managers trade indicates that they don’t value picks by expected performance or surplus value. In particular, general managers trade suboptimally: trading down in the draft can be incredibly valuable, and they often overpay when they trade up.
As shown in Section 1.2, Massey and Thaler’s analysis from over a decade ago completely replicates to recent data (2013-2023). Even though their work has been public and widely known for a long time, general managers’ draft behavior has remained largely unchanged. One explanation is that general managers are stubborn and haven’t learned from these analyses. If you believe general managers should value picks by expected performance or surplus value, they have continued to make terrible trades on average, not trading down often enough and overpaying when they trade up. An alternative explanation is that there is a sound economic reason why average future player value (at a draft position) doesn’t determine trade value. General managers are using some other valuation function that captures an essential piece of the puzzle missing from previous analyses. We are partial to the latter explanation.
A general manager’s goal is to win the Super Bowl. Winning the Super Bowl entails being the best team of 32. Mathematically, thinking of each team’s performance as a random variable, being the best team is like exceeding a high threshold of team performance. Having the highest expected value is not necessarily the best way to accomplish this. If high expected value comes with little variance, you may never exceed a high threshold.
Consider a toy example where the performance of each football team is determined by a Guassian random variable. Suppose team $A$’s performance has mean 1 and s.d. 0 and team $B$’s has mean 0 and s.d. 1. Suppose also that to have a good shot at winning the Super Bowl, a team needs to exceed a performance value of $1.28$. Though team $A$’s performance is much higher than $B$’s on average, $A$’s probability of exceeding the threshold is 0 and $B$’s is 10%. Team $B$ has a decent shot because its high variance puts nontrivial probability mass on far-right tail performance.
With this in mind, we consider how variance of performance changes as the draft progresses. In Figure 2 we visualize the conditional mean $x\mapsto{\mathbb{E}}[Y|x]$ and standard deviation $x\mapsto{\mathrm{sd}}(Y|x)$ of performance $Y$ as a function of draft position $x$. These graphs look remarkably similar. The key insight is that the conditional mean and variance of performance $Y$ decreases as draft pick $x$ increases. Moreover, they decrease convexly. Since an earlier pick has a much higher expected value and variance than a later pick, it has a much fatter right tail. In other words, an earlier pick has a significantly higher probability of becoming an elite player, moreso than we would expect if we just considered expected value.
@@ 1.3 Research hypothesis
Figure 2: The black dots indicate the empirical mean (left) and s.d. (right) of $Y$ (performance, i.e. second contract APY cap percentage) given $x$ (draft position). The blue lines are smoothed curves $x\mapsto{\mathbb{E}}[Y|x]$ (left) and $x\mapsto{\mathrm{sd}}(Y|x)$ (right).
If a general manager’s goal is to win a Super Bowl, then accepting a trade because it yields higher performance or surplus value on average is not necessarily optimal. We suspect general managers also value variance, which puts subtantial probability mass on far-right tail outcomes. In other words, we suspect general managers value performance nonlinearly, placing exponentially higher value on players as their eliteness increases. Intuitively, elite players in the far-right tail of performance are essential to a team’s performance exceeding a high threshold.
Ideally we would quantify the probability of winning a Super Bowl if a team drafts a player of performance value $Y$, but this is extremely complex for many reasons and we leave it to future work (see Section 3 for details). Instead, in this paper we consider simpler alternative value functions beyond expected performance or surplus value. We focus on nonlinear transformations of $Y$ that capture the outsize influence of elite players. The most intuitive example is right tail probability, which places high value on elite performance and no value on lesser performance. We find that such nonlinear value functions produce steeper draft value curves that more closely resemble the observed trade market. If you are looking for elite talent, picks from lower draft picks aren’t worth much and the high cost of a top draft pick is justified. Additionally, we adjust for position. In particular, we build draft value curves for quarterbacks and non-quarterbacks. We find that, if you are looking for an elite quarterback, there is no loser’s curse. The far-right tails of performance $Y$ are so fat for top pick quarterbacks that even the high cost of top picks doesn’t massively diminish their value.
The remainder of this paper is organized as follows. In Section 2 we build draft value curves that value eliteness and variance. We begin with right tail probability and extend to more intricate valuations. We then consider surplus value and adjust for position. In Section 3 we conclude.
2 Draft curves that value eliteness
@@ 2.1 Right tail probability
An intuitive example of valuing a draft position by eliteness is right tail probability. Formally, we consider the probability that the performance $Y$ of a player drafted at position $x$ exceeds a threshold $r$, ${\mathbb{P}}(Y>r|x)$. The tail cutoff $r$ is defined by a percentage of the salary cap, not a quantile. For example, a general manager may be interested in the probability that a drafted player is at least as valuable as Cowboys’ quarterback Dak Prescott ($r=15.8\%$). Since the expected value of an indicator is its probability, this is equivalent to the number of elite players expected to arise from a given pick.
We compute right tail probability from the conditional density ${\mathbb{P}}(Y|x)$, which we estimate using a Bayesian Beta regression model, detailed in Appendix D.1. As both the mean and variance of $Y$ decay convexly over the draft (recall Figure 2), the right tail decays convexly as well, which we visualize in Figure 3. The densities of top picks feature fat right tails, which shift lefward and morph into a spike near 0 as the draft progresses (see Figures 12 and 14 in Appendix D.1).
@@
Figure 3: Right tail probability ${\mathbb{P}}(Y>r|x)$ ($y$-axis) as a function of draft position $x$ and eliteness cutoff $r$ (color).
In Figure 4 we visualize right tail probability draft value curves. We overlay the expected performance value curve (from Figure 1) and the fitted trade market curve (from Appendix C). The right tail curves are steeper than the expected value curve. To draft for eliteness, not expected value, earlier picks are more valuable relative to later picks. Futher, as the eliteness cutoff increases, the right tail draft value curve gets steeper and steeper until it nearly matches the fitted trade market curve. In particular, general managers trade draft picks as if they value a draft position by the probability it results in a player whose second contract exceeds $15\%$ of the salary cap. Though we can’t peer into the minds of general managers, if they were trading for eliteness, we would have seen a similar trade market. This is a high bar of eliteness: in our dataset of draft picks from 2013 to 2023, just 16 players ($\approx 0.57\%$ of players) satisfy $Y\geq 15\%$.
@@
Figure 4: Performance value relative to the first pick ($y$-axis) as a function of draft position $x$. The colored lines are right tail probability curves proportional to ${\mathbb{P}}(Y>r|x)$, the dotted black line is proportional to expected performance value ${\mathbb{E}}[Y|x]$, and the solid black line is the fitted trade market curve.
2.2 S curves
[…]
Figure 5: Performance outcome value functions ${g_{\text{step}}}(y)=\mathbb{I}\left(y>r\right)$ (a), two ${g_{\text{s}}}$ curves (b,c), and ${g_{\text{line}}}(y)=y$ (d).
An oddity of the step function ${g_{\text{step}}}$ is that it so sharply characterizes success and failure: $Y=r$ is a total failure and $Y=r+\varepsilon$ is a total success. A weakness of the linear function ${g_{\text{line}}}$ is that it discounts the outsized influence of elite players on winning. For instance, it values the difference between $2\%$ and $1\%$ the same as the difference between $9\%$ and $8\%$. If you believe great players are exponentially more important than typical players for winning games, it is critical to use a nonlinear $g$.
Hence, we propose a nonlinear transformation $g$ between the two extremes of ${g_{\text{step}}}$ and ${g_{\text{line}}}$. To nonlinearly transform performance but in a less sharp way, we use an $s$ curve. We use the CDF of the beta distribution, a sufficiently rich family of $s$ curves, denoted by ${g_{\text{s}}}$ and parameterized by $\alpha$ and $\beta$,
[…]
We visualize two of these $s$ curves in Figures 5b and 5c. For example, ${g_{\text{s}}}$ in Figure 5b assigns near failure to busts and near total success to elite players. The two dashed gray lines indicate where ${g^{\prime}_{\text{s}}}(y)=1$. When $y$ is between the two gray lines, ${g^{\prime}_{\text{s}}}(y)>1$. This reflects, for example, that the difference in value between a good and median player is much larger for ${g_{\text{s}}}$ than for ${g_{\text{line}}}$ (since ${g_{\text{line}}}^{\prime}(y)\equiv 1)$. Similarly, the difference in value between a median and a bad player is much larger for ${g_{\text{s}}}$ than for ${g_{\text{line}}}$.
An $s$ curve ${g_{\text{s}}}$ produces a draft value curve ${v_{\text{s}}}(x)\propto{\mathbb{E}}[{g_{\text{s}}}(Y)|x]$, which we visualize in Figure 6. The $s$ curve draft curves ${v_{\text{s}}}$ (green and purple dashed lines) lie between expected performance (black dotted line) and right tail probability curves (red and blue solid lines). The ${v_{\text{s}}}$ curves lie between the two extremes because they place exponentially increasing value on increasingly good players while still placing nonzero value on less-than-elite production.
Figure 6: The value $v(x)\propto{\mathbb{E}}[g(Y)|x]$ of draft position $x$ relative to the first pick ($y$-axis) for various transformations $g$ of performance (color). Draft curves arising from step function transformations $g={g_{\text{step}}}$ are solid lines, from $s$ curve transformations $g={g_{\text{s}}}$ are dashed lines, and from the linear function $g={g_{\text{line}}}$ are dotted lines.

### v2

@@ 3 Traditional draft position value curves
The trade market curve (solid black line) is much steeper than the expected performance value curve (dotted black line) and the expected surplus value curve (long-dashed black line). Teams aren’t valuing draft positions by expected performance or surplus value. The implication is that general managers are trading suboptimally: they don’t trade down often enough, and they often overpay when they trade up. According to expected surplus value, even teams at the very top of the draft should trade down, particularly towards the middle of the first round. This result is counterintuitive because very top picks are so highly regarded by general managers and football fans.
4 Research hypothesis
Massey and Thaler, (2013) pioneered the traditional analytical approach to constructing draft position value curves. Their study produced two key findings. First, they discovered the loser’s curse––last years worst teams are “cursed” with the ostensible gifts of the top picks, which cost a lot (in compensation) without delivering enough on-field value to compensate. Second, they demonstrated that the draft pick trade market is highly inefficient––teams significantly overvlaue top draft picks, consistently overpaying to trade up in the draft without trading down often enough.
Despite the widespread dissemination of these findings, general managers’ draft behavior has remained largely unchanged. When applied to recent data (2013––2023), the loser’s curse persists and the fitted trade market curve remains similar (see Section 3 and Avery et al., ming, forthcoming). If general managers were optimizing for expected performance or surplus value, their continued reluctance to trade down and their willingness to pay a premium for higher picks would suggest systematic misjudgment. One possible explanation is that general managers have simply failed to learn from these analyses. An alternative explanation is that general managers are instead operating under a fundamentally different utility function. If maximizing expected surplus value does not align with their actual objectives, then their observed behavior may be rational, even if it appears suboptimal under Massey and Thaler’s framework. This paper challenges their core assumption, that general managers should adopt expected surplus value as their utility function in valuing draft picks. This assumption is neither explicitly justified in their analysis nor necessarily optimal in the context of NFL decision making.
A general manager’s primary objective is to construct a Super Bowl-winning roster. Achieving this goal requires assembling a team capable of reaching and surpassing an elite performance threshold. Mathematically, thinking of team performance as a random variable, winning the Super Bowl is analagous to exceeding a high threshold of team performance relative to other teams. In such a setting, maximizing expected value alone may not be optimal. A strategy that emphasizes high-variance outcomes—seeking players with the potential for extreme right-tail performance—may be preferable, even if it sacrifices consistency.
To illustrate this concept, consider a simple example in which team performance follows a Gaussian distribution. Suppose team $A$ has performance distributed as ${\mathcal{N}}(1,0)$ (mean 1, variance 0), while team $B$’s performance follows ${\mathcal{N}}(0,1)$ (mean 0, variance 1). If winning the Super Bowl requires exceeding a performance threshold of 1.28, team $A$’s probability of success is zero, whereas team $B$’s is $1/10$, despite having a lower expected value. This example highlights how variance, not just expectation, plays a crucial role in achieving extreme outcomes.
This framework has direct implications for evaluating draft picks. If elite players––those in the far right tail of the performance distribution––are disproportionately valuable in constructing a championship team, then upweighting variance in draft selections is rational. Figure 3 shows that both the expected value and variance of performance decreases as draft position increases. Critically, the decline in variance suggests that higher draft picks offer significantly greater potential for elite outcomes, reinforcing the rationale for teams to value them more than suggested by expected surplus value models.
@@ 4 Research hypothesis
Figure 3: The gray dots denote the empirical mean (left) and s.d. (right) of performance $Y$ given draft position $x$. The black lines are smoothed curves $x\mapsto{\mathbb{E}}[Y|x]$ (left) and $x\mapsto{\mathrm{sd}}(Y|x)$ (right).
Thus, rather than viewing general managers’ draft strategies as systematically flawed, we propose that their behavior aligns with a decision rule that accounts for variance and right-tail outcomes. Our hypothesis is that, in trading draft picks, teams prioritize something closer to the probability of drafting transformational players over maximizing expected surplus value. These alternative valuation functions lead to steeper draft value curves that are more consistent with observed trade behavior.
5 Conditional density estimation
By modeling the distribution of performance as a function of relevant covariates, we can construct a broad class of draft position value curves that extend beyond expected performance or surplus value. Hence, in this section, we estimate the conditional density of performance given draft position (Section 5.1) and given both draft position and player position (Section 5.2).
5.1 Position-agnostic conditional density
In Figure 4 we visualize the empirical conditional density of performance $Y$ given draft position $x$. The empirical densities are extremely wobbly and noisy due to limited data for each individual draft position. But there is a general trend: we see a spike near zero that grows upwards as draft position increases and a right tail that shifts leftwards and diminishes as draft position increases. This motivates our “Bayesian spike plus Beta regression” model of ${\mathbb{P}}(Y|x)$.101010 We initially tried a lone Beta regression model without the spike near zero, which less successfully fit the right tail.
Figure 4: Empirical distribution (histogram and smoothed density) of performance $Y$ ($x$-axis) given draft position $x$ (facet).
We model the spike near zero, the bust spike, using logistic regression. We let a bust be a player whose performance outcome $Y$ is smaller than some cutoff ${y_{\text{bust}}}$. We found that ${y_{\text{bust}}}=0.5\%$ of the cap works well. We model the bust probability ${\mathrm{bp}}(x):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$ by
$1\{Y\leq{y_{\text{bust}}}\}|x\sim{\mathrm{Bernoulli}}({\mathrm{bp}}(x)),$ (5.1)
where
${\mathrm{bp}}(x)={\mathrm{logistic}}(\alpha_{0}+\alpha_{1}\cdot x).$ (5.2)
Next, we model the right tail using Beta regression. We view the Beta distribution as a natural way to model $Y$ because it lies in $[0,1]$ and aligns well with the empirical densities. We model
$Y|x,Y>{y_{\text{bust}}}\sim{\mathrm{Beta}}(\mu(x),\phi(x)).$ (5.3)
Here, we use the mean-precision parameterization of the Beta distribution. The mean $\mu$ and precision $\phi$ are related to the traditional Beta distribution parameters $\mathrm{shape}_{1}$ and $\mathrm{shape}_{2}$ by $\mathrm{shape}_{1}=\mu\cdot\phi$ and $\mathrm{shape}_{2}=(1-\mu)\cdot\phi$. We let the mean and precision parameters vary as draft position $x$ varies via $\mu(x)$ and $\phi(x)$ to capture that different draft positions $x$ yield different outcome distributions. The conditional mean is ${\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]=\mu(x)$ and the conditional variance is $\mathrm{Var}[Y|x,Y>{y_{\text{bust}}}]=\mu(x)\cdot(1-\mu(x))/(1+\phi(x))$. So, the precision is inversely proportional to the variance. We use spline regression to model the conditional mean,
$\mu(x)={\mathrm{logistic}}(\widetilde{x}^{\top}\beta),$ (5.4)
where $\widetilde{x}$ is a B-spline basis.111111 The ${\mathsf{R}}$ code we used to generate the spline basis is splines::bs($x$, degree=3, df=4, intercept=T). We model the conditional precision by
$\phi(x)=\exp(\gamma_{0}+\gamma_{1}\cdot x).$ (5.5)
To capture uncertainty in the parameters, we use a Bayesian model, placing priors on the parameters. We put a weakly informative Normal prior with mean zero and s.d. ten on each of the $\alpha$, $\beta$, and $\gamma$ parameters. Finally, we assume that $Y$ is uniformly distributed within the bust spike in a way that doesn’t vary with $x$: $Y|x,Y\leq{y_{\text{bust}}}\sim{\mathrm{Unif}}[0,{y_{\text{bust}}}]$.121212 We assess the reasonableness of this assumption later in this section.
In Figure 5 we see that our estimates of the conditional mean $\mu(x)$, conditional standard deviation ${\mathrm{sd}}(x)$, and bust probability ${\mathrm{bp}}(x)$ fit the data well. In Figure 6 we see that our estimates of the right tail ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ fit the data well enough. As expected, the right tail shifts leftwards as draft position increases. In Figure 7 we visualize the conditional density ${\mathbb{P}}(Y|x)$ for values $Y$ in the right tail $Y>{y_{\text{bust}}}$. By Bayes rule, this is equal to ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$. The right tail, measuring the probability of drafting an elite player, shifts leftwards and decreases in size as draft position increases. In Figure 8 we visualize the conditional density of surplus $S=Y-{\mathrm{cost}}(x)$. The high cost of first round picks shifts those curves so much to the left that there is substantial probability mass on negative (below-replacement) outcomes.
Figure 5: Left: our estimate of the conditional mean $x\mapsto\mu(x)={\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]$ given not a bust. Middle: our estimate of the conditional standard deviation $x\mapsto{\mathrm{sd}}(x)={\mathrm{sd}}(Y|x,Y>{y_{\text{bust}}})$ given not a bust. Right: our estimate of bust probability $x\mapsto{\mathrm{bp}}(x)={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$. The solid black lines are the posterior means, the cyan shaded regions are the $95\%$ credible intervals, and the gray dots are the empirical mean values for each draft position $x$.
Figure 6: The distribution of performance $Y$ ($x$-axis) given draft position $x$ (facet). The solid black lines are the fitted conditional densities ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ (the posterior mean curves) and the cyan shaded regions are the $95\%$ credible intervals, which overlay the empirical distributions (histogDolphins and smoothed densities) in gray.
Figure 7: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional density ${\mathbb{P}}(Y|x)={\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$ for various values of draft position $x$ (color). The lines are posterior mean density curves and the shaded regions are 95% credible intervals of those curves.
Figure 8: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional surplus density ${\mathbb{P}}(S|x)={\mathbb{P}}(Y-{\mathrm{cost}}(x)|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$ for various values of draft position $x$ (color) in the first round (a) and throughout the draft (b). The lines are posterior mean density curves and the shaded regions are 95% credible intervals of those curves.
Figure 9: The gray dots display the empirical mean (left) and standard deviation (right) of performance $Y$ given draft position $x$ conditional on being a bust $Y\leq{y_{\text{bust}}}$. The black lines are smoothed lines.
5.2 Position-specific conditional density
Now, we extend the “Bayesian spike plus Beta regression” model from the previous section to adjust for player position, henceforth referred to as just “position” and denoted ${\mathsf{pos}}$. We use 11 position groups: quarterback (QB), wide receiver (WR), running back / fullback (RB/FB), tight end (TE), offensive tackle (OT), interior offensive lineman (IOL), interior defensive lineman (DL), edge (ED), linebacker (LB), cornerback (CB), and safety (S). We filter out all draft picks from our dataset belonging to other positions (e.g., punter, kicker, etc.).
It is imperative to use a Bayesian model to adjust for position, rather than fitting separate curves for each position, due to limited data. Across each of the positions, the median sample size in our dataset is 269 players, with a maximum of 350 and a minimum of 127. Spread across 256 total draft pick positions, we are left with an average of about one observation per combination of position and draft position. A Bayesian hierarchical model is an excellent choice here because it allows us to share information across observations via shrinkage and to capture uncertainty in position curves.
$Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}}\sim{\mathrm{Beta}}(\mu(x,{\mathsf{pos}}),\phi(x,{\mathsf{pos}})),$ (5.6)
where
$\mu(x,{\mathsf{pos}})={\mathrm{logistic}}(\widetilde{x}^{\top}\beta_{\mathsf{pos}})$ (5.7)
and
$\phi(x,{\mathsf{pos}})=\exp(\gamma_{0,{\mathsf{pos}}}+\gamma_{1,{\mathsf{pos}}}\cdot x).$ (5.8)
${\mathrm{bp}}(x,{\mathsf{pos}}):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x,{\mathsf{pos}})={\mathrm{logistic}}(\alpha_{0,{\mathsf{pos}}}+\alpha_{1,{\mathsf{pos}}}\cdot x).$ (5.9)
We shrink the position-specific coefficients $\alpha_{\ast,{\mathsf{pos}}}$, $\beta_{\ast,{\mathsf{pos}}}$, and $\gamma_{\ast,{\mathsf{pos}}}$ towards overall mean parameters $\alpha_{\ast}$, $\beta_{\ast}$, and $\gamma_{\ast}$. For example, we model the position-specific coefficients $\alpha_{0,{\mathsf{pos}}}$ as a draw from a Normally distributed prior with mean $\alpha_{0}$ and standard deviation $\tau_{\alpha_{0}}$. We put a weakly informative Normal prior with mean zero and s.d. ten on the prior mean parameter $\alpha_{0}$. We put a mild positive standard Normal prior on the prior standard deviation parameter $\tau_{\alpha_{0}}$. The priors and hyperpriors are the same for the other parameters.
In Figure 10a we see that our estimate of the conditional mean $\mu(x,{\mathsf{pos}})$ fits the data well. The empirical standard deviation values in Figure 10b we are much more sparse than the empirical conditional mean values in Figure 10a. This is because, for each position ${\mathsf{pos}}$ and for many draft positions $x$, just zero or one player of position ${\mathsf{pos}}$ was drafted at $x$. Thus, there is no empirical standard deviation at many combination of $x$ and ${\mathsf{pos}}$. Nonetheless, based on the data we have observed, our estimate of the conditional standard deviation ${\mathrm{sd}}(x,{\mathsf{pos}})$ fits the data reasonably well.
Figure 10: Our estimates of the conditional mean $\mu(x,{\mathsf{pos}})={\mathbb{E}}[Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}}]$ (a) and standard deviation ${\mathrm{sd}}(x,{\mathsf{pos}})={\mathrm{sd}}(Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}})$ (b) given not a bust. The solid black lines are the posterior means, the cyan shaded regions are the $95\%$ credible intervals, and the gray dots are the empirical base-rates.
In Figure 11a we visualize our estimates of the conditional mean, s.d., and bust probability given position and draft position. Though the posterior mean curves vary by position, the posterior credible intervals are wide. The non-quarterback mean and s.d. curves jumble into one big clump. Hence, in Figure 11b we average all of the non-quarterback curves into the black curves, $\mu(x,{\mathsf{not\ qb}})=\frac{1}{\#{\mathsf{pos}}-1}\sum_{{\mathsf{pos}}^{\prime}\neq{\mathsf{qb}}}\mu(x,{\mathsf{pos}}^{\prime})$ and similarly for ${\mathrm{sd}}(x,{\mathsf{not\ qb}})$ and ${\mathrm{bp}}(x,{\mathsf{not\ qb}})$. As expected, quarterbacks have a higher mean than other positions. More interestingly, quarterbacks have a much higher variance than other positions. This is what drives the fat far right tail of top-pick quarterback performance.
Figure 11: Figure (a): The posterior mean estimates (lines) and 95% credible intervals (shaded regions) of $\mu(x,{\mathsf{pos}})$ (left), ${\mathrm{sd}}(x,{\mathsf{pos}})$ (middle), and ${\mathrm{bp}}(x,{\mathsf{pos}})$ (right) as a function of draft position $x$ and position ${\mathsf{pos}}$ (color). Figure (b): These curves for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$.
Figure 12: For values of $Y$ in the right tail $Y>{y_{\text{bust}}}$, the posterior mean estimates (lines) and 95% credible intervals (shaded regions) of the conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})={\mathbb{P}}(Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x,{\mathsf{pos}}))$, for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (facet) and various values of draft position $x$ (color) throughout the draft.
In Figure 12 we visualize the conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})$ for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (where the ${\mathsf{not\ qb}}$ density is the average of all the non-quarterback positional densities). As the draft progresses, for both ${\mathsf{qb}}$’s and ${\mathsf{not\ qb}}$’s the right tail shifts leftwards and decreases in size as draft position increases. Notably, the ${\mathsf{qb}}$ right tails are much fatter than the ${\mathsf{not\ qb}}$ tails. This reflects the immense value of quarterbacks, and top-pick quarterbacks in particular.
6 Results
Given our estimated conditional density of performance given draft position and player position, in this section we construct draft position value curves from alternative objective functions. We move beyond expected performance and surplus value to objective functions that emphasize elite, right-tail outcomes.
@@ 6.1 Right tail probability
An intuitive approach to valuing a draft position in a way that prioritizes eliteness is by right-tail probability. Formally, we consider the probability that the performance $Y$ of a player drafted at position $x$ exceeds a threshold $r$, ${\mathbb{P}}(Y>r|x)$. The threshold $r$ is determined relative to a percentage of the NFL salary cap. For example, a general manager may be interested in the probability that a drafted player achieves at least the performance level of quarterback Derek Carr ($r=15.0\%$). Since the expected value of an indicator function is its probability, this equivalently measures the number of elite players expected to arise from a given pick.
We compute right tail probabilities from the estimated conditional density ${\mathbb{P}}(Y|x)$. As both the mean and variance of $Y$ decrease convexly as the draft progresses (see Figure 3)––because the densities of top picks feature fat right tails, which shift lefward and morph into a spike near zero as the draft progresses (see Figure 7)––the right tail of the performance distribution exhibits a similar convex decay, which we visualize in Figure 13. We consider five eliteness cutoffs $r$: $24.5\%$, $17.8\%$ $15.0\%$, $11.0\%$ and $8.7\%$. Among quarterbacks in our dataset who secured multi-year second contracts as regular starters, Joe Burrow has the highest performance value ($r=24.5\%$) and Derek Carr has the lowest ($r=15.0\%$). Jared Goff lies near the midpoint of these two benchmarks ($r=17.8\%$). Among non-quarterbacks, offensive tackle Penei Sewell ranks at the $99^{th}$ percentile ($r=11.0\%$), while cornerback Tre’Davious White lies at the $90^{th}$ percentile ($r=8.7\%$).
@@
Figure 13: Estimated right tail probability ${\mathbb{P}}(Y>r|x)$ ($y$-axis) as a function of draft position $x$ and eliteness cutoff $r$ (color). The lines are the posterior means and the shaded regions are the $95\%$ credible intervals.
@@
Figure 14: Value relative to the first pick ($y$-axis) as a function of draft position $x$. The dotted black line is the expected performance value curve. The solid black line is the fitted trade market curve. The colored lines are proportional to right tail probability $x\mapsto{\mathbb{P}}(Y>r|x)$, where $r$ is an eliteness cutoff (color). The lines are the posterior means and the shaded regions are the $95\%$ credible intervals.
In Figure 14 we visualize draft position value curves proportional to right tail probability, $x\mapsto{\mathbb{P}}(Y>r|x)/{\mathbb{P}}(Y>r|x=1)$. We overlay the expected performance value curve and the fitted trade market curve from Figure 2. Notably, the right tail probability curves are steeper than the expected performance value curve. This suggests that if teams prioritize eliteness rather than expected performance, earlier draft picks hold substantially greater value relative to later picks. Furthermore, as the eliteness threshold $r$ increases, right-tail draft value curves progressively steepen, ultimately converging to or even surpassing the fitted trade market curve.
Further, we find that general managers trade draft picks as if they value a draft position by the probability it produces a player whose second contract exceeds $19.7\%$ of the salary cap. Specifically, the mean absolute error between the fitted trade market curve and a right tail probability curve with eliteness threshold $r$ is minimized by $r=19.7\%$ (note that Deshaun Watson has $r=19.7\%$). This is a high bar of eliteness: among quarterbacks in our dataset who secured multi-year second contracts as regular starters, just $12.3\%$ of them ($8/65$) met this criterion. Although we cannot directly observe general managers’ internal valuation models, the trade market aligns closely with what we would expect if teams were optimizing for elite player acquisition rather than expected performance.